The following is a partial trial balance for the Green Star Corporation as of December 31, 2018:

Account Title Debits Credits
Sales revenue 1,850,000
Interest revenue 44,000
Gain on sale of investments 64,000
Cost of goods sold 830,000
Selling expenses 230,000
General and administrative expenses 89,000
Interest expense 54,000
Income tax expense 144,000


130,000 shares of common stock were outstanding throughout 2018.

Required:
1. Prepare a single-step income statement for 2018, including EPS disclosures.
2. Prepare a multiple-step income statement for 2018, including EPS disclosures.
Answer #1
Solution:
1.
GREEN STAR CORPORATION
Income Statement
For the Year Ended December 31, 2018
Revenues and gains:
a Sales revenue 1,850,000
b Interest revenue 44,000
c Gain on sale of investments 64,000
d=a+b+c Total revenues and gains 1,958,000
Expenses and losses:
e Cost of goods sold 830,000
f Selling expenses 230,000
g General and administrative expenses 89,000
h Interest expense 54,000
i=e+f+g+h Total expenses and losses 1,203,000
j=d-i Income before income taxes 755,000
k Income tax expense 144,000
l=j-k Net income 611,000
m=l/130,000 Earnings per share 4.70
Notes: Earnings per share= Net income/shares of common stock were outstanding
=611,000/130,000 = 4.70
2.
GREEN STAR CORPORATION
Income Statement
For the Year Ended December 31, 2018
a Sales revenue 1,850,000
b Cost of goods sold (830,000)
c=a-b Gross profit 1,020,000
Operating expenses:
d Selling expenses (230,000)
e General and administrative expenses (89,000)
f=d+e Total operating expenses (319,000)
g=c-f Operating income 701,000
Other income (expense):
h Interest revenue 44,000
i Gain on sale of investments 64,000
j Interest expense (54,000)
k=h+i-j Total other income, net 54,000
l=g+k Income before income taxes 755,000
m Income tax expense (144,000)
n=l-m Net income 611,000
o=n/130,000 Earnings per share 4.70
Notes: Earnings per share= Net income/shares of common stock were outstanding
=611,000/130,000 = 4.70
Please feel free to ask if anything about above solution in comment section of the question.